OpenStudy (anonymous):

Given the following roots determine the polynomial: -4(multiplicity of 2), 3

OpenStudy (amoodarya):

(x-3) (x-(-4))(x-(-4))

OpenStudy (anonymous):

(x−3)(x−(−4))(x−(−4)) =((x−3)(x−(−4)))(x+4) =((x−3)(x−(−4)))(x)+((x−3)(x−(−4)))(4) =−12x+x2+x3+4x+4x2−48 =x3+5x2−8x−48

OpenStudy (anonymous):

You must break it down @amoodarya Or else your equations are useless in helping